# Gapfree graphs and powers of edge ideals with linear quotients

### v1

@@ 4. Pentagon together with vertices
The edge ideal $I(G)$ has $8$ generators and, since $e_{2}e_{6}=e_{4}e_{7}$ and $e_{3}e_{8}=e_{5}e_{7}$, it follows that $I(G)^{2}$ has $\binom{8+1}{2}-2=36-2=34$ generators. Let ${\mathcal{N}}$ denote the following order these generators:
@@
It is now a direct computation to verify that with the order ${\mathcal{N}}$, $I(G)^{2}$ has linear quotients.
Using the ordering ${\mathcal{N}}$ of the generators of $I(G)^{2}$ as above, one can inductively show that all powers of $I(G)$ have linear quotients. The following notation will be useful in the proof of Proposition 4.2.
Definition 4.1.
Let $G$ be a finite graph and $w$ one of the generators of $I(G)^{t}$. The notation $e_{j}\shortmid w$ means that there exists a factorization $w=e_{i_{1}}\cdots e_{i_{t}}$ with $e_{j}=e_{i_{k}}$ for some $1\leq k\leq t$.
@@
Let $G$ be the finite graph of Figure $2$. Then all powers $I(G)^{N}$ with $N\geq 2$ have linear quotients.
@@
We prove that Conjecture 3.1 is true for $G$ with $q_{0}=2$. Assume $N\geq 3$ and suppose that $I(G)^{N-1}$ has linear quotients with the ordering ${\mathcal{N}}$: $u_{1}>\cdots>u_{r}$ where $u_{1}=e_{1}^{N-1}$ and $u_{r}=e_{8}^{N-1}$. Let ${\mathcal{N}}^{(1)}:~{}~{}w_{1}>\cdots>w_{t}$ denote the efficient ordering of $I(G)^{N}$ coming from ${\mathcal{N}}$ as in Conjecture 3.1. For ease of notation, use the following notation to refer to elements in ${\mathcal{N}}^{(1)}$
${\mathcal{U}}e_{1}={\mathcal{A}}(e_{1}),{\mathcal{A}}(e_{4}),{\mathcal{A}}(e_{2}),{\mathcal{A}}(e_{3}),{\mathcal{A}}(e_{5}),{\mathcal{A}}(e_{6}),{\mathcal{A}}(e_{7}),{\mathcal{A}}(e_{8})$
where $w_{k}\in{\mathcal{A}}(e_{i})$ means $e_{i}\shortmid w_{k}$ and $e_{j}\nshortmid w_{k}$ for each $j<i$.
@@
Now, let $w_{n}=u_{s}e_{4}\in{\mathcal{A}}(e_{4})$ for $s\leq r$. First, observe that $(b)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:u_{s}e_{4})=(b)$ for $u_{s}e_{1}\in{\mathcal{A}}(e_{1})$. Let $w_{i}=u_{t}e_{4}\in{\mathcal{A}}(e_{4})$ for $t<s$. By the induction hypothesis, there exists $u_{l}$ with $l<s$ such that $u_{l}:u_{s}$ is a variable dividing $u_{t}:u_{s}$. Let $w_{j}=u_{l}e_{4}$ for $j<n$. Then, $w_{j}:w_{n}$ is a variable and it divides $w_{i}:w_{n}$, as desired. Note that $w_{j}\in{\mathcal{A}}(e_{1})$ is possible.
Next, let $w_{i}\in{\mathcal{A}}(e_{1})$ and consider whether $b$ divides $w_{n}=u_{s}e_{4}$. Recall that $b\nmid e_{4}$. If $b\nmid u_{s}$, then $w_{i}:w_{n}\in(b)$ since every element of ${\mathcal{A}}(e_{1})$ is divisible by $b$. Then, $(u_{1}e_{1},\ldots,u_{r}e_{1}):w_{n}=(b)$. Hence, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case. Suppose $b\mid u_{s}$. Then either $e_{3}\shortmid u_{s}$ or $e_{5}\shortmid u_{s}$. If $e_{3}\shortmid u_{s}$, then $u_{s}=w^{\prime}e_{3}$ for a generator $w^{\prime}\in I(G)^{N-2}$. Consider the following elements $w_{j}=w^{\prime}e_{4}e_{1}$ and $w_{k}=w^{\prime}e_{7}e_{1}$ in ${\mathcal{A}}(e_{1})$ where $(w_{j}:w_{n})=(a)$ and $(w_{k}:w_{n})=(z)$. If $e_{5}\shortmid u_{s}$, then $u_{s}=w^{\prime\prime}e_{5}$ for a generators $w^{\prime\prime}\in I(G)^{N-2}$. Similarly, consider $w_{j}=w^{\prime\prime}e_{4}e_{1}$ and $w_{k}=w^{\prime\prime}e_{8}e_{1}$ in ${\mathcal{A}}(e_{1})$ where $(w_{j}:w_{n})=(a)$ and $(w_{k}:w_{n})=(z)$. So, in either of these scenarios, we have $(a,b,z)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}.$ Let $w_{i}=e_{1}^{t_{1}}\cdots e_{8}^{t_{8}}$ where $t_{1}>0$ and $w_{n}=e_{2}^{t_{2}^{\prime}}\cdots e_{8}^{t_{8}^{\prime}}$ since $e_{1}\nshortmid w_{n}$. If $w_{i}:w_{n}\notin(a,b,z)$, then examining the degrees of $a,b,z$ in $w_{i}$ and $w_{n}$, respectively, results with
@@
Let $w_{n}=u_{s}e_{2}\in{\mathcal{A}}(e_{2})$ for some $s\leq r$. Note that $e_{6}\nshortmid w_{n}$ because $e_{6}e_{2}=e_{7}e_{4}$ and $e_{4}\nshortmid e_{n}$. Next, observe that $(b,p)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:u_{s}e_{2})=(b)$ and $(u_{s}e_{4}:u_{s}e_{2})=(p)$. Let $w_{i}\in{\mathcal{A}}(e_{2})$. As in the previous case, by the induction hypothesis, there exists $w_{j}$ for $j<n$ such that $w_{j}:w_{n}$ is a variable and it divides $w_{i}:w_{n}$, as desired. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})$. Observe that $p\nmid u_{s}$; otherwise, $e_{4}\shortmid u_{s}$ or $e_{6}\shortmid u_{s}$ but neither of these cases are possible since $w_{n}\in{\mathcal{A}}(e_{2})$. If $b\nmid u_{s}$, then $(w_{i}:w_{n})\subseteq(b,p)$ which implies that $(w_{1},\ldots,w_{n-1}):w_{n}$ is generated by variables. So, claim holds for this case. Suppose $b\mid u_{s}.$ Then either $e_{3}\shortmid u_{s}$ or $e_{5}\shortmid u_{s}$ since $e_{1}\nshortmid u_{s}.$ If $e_{3}\shortmid u_{s}$, then $u_{s}=w^{\prime}e_{3}=w^{\prime}bx$ for some $w^{\prime}\in I(G)^{N-2}$. Consider the elements $w_{j}=w^{\prime}e_{2}e_{1}$ and $w_{k}=w^{\prime}e_{7}e_{1}$ in ${\mathcal{A}}(e_{1})$ where $(w_{j}:w_{n})=(a)$ and $(w_{k}:w_{n})=(z)$. So, $(a,b,p,z)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}.$
@@
Assuming $e_{5}\shortmid u_{s}$ also yields to the above containment. Let $w_{i}=e_{1}^{t_{1}}\cdots e_{8}^{t_{8}}$ where $t_{1}>0$ or $t_{4}>0$ and $w_{n}=e_{2}^{t_{2}^{\prime}}e_{3}^{t_{3}^{\prime}}e_{5}^{t_{5}^{\prime}}e_{t}^{t_{7}^{\prime}}e_{8}^{t_{8}^{\prime}}$ since $e_{1},e_{4},e_{6}$ do not divide $w_{n}=u_{s}e_{2}.$ If $w_{i}:w_{n}\notin(a,b,p,z)$, then examining degrees of $a,b,p,z$ in $w_{i}$ and $w_{n}$ results with $t_{1}+t_{4}+t_{6}\leq 0$, a contradiction. So, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
@@
Now, let $w_{n}=u_{s}e_{3}\in{\mathcal{A}}(e_{3})$ for some $u_{s}\in I(G)^{N-1}$. First observe that $(a)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:u_{s}e_{3})=(a)$. The case $w_{i}\in{\mathcal{A}}(e_{3})$ follows from the induction hypothesis as in the previous cases. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})\cup{\mathcal{A}}(e_{2})$. Note that this collection contains all generators of $I(G)^{N}$ that are divisible by $a$ since $e_{1},e_{2}$ and $e_{4}$ are the only edges incident to $a$ and $a\nmid w_{n}$. This means $w_{i}:w_{n}\in(a)$. Hence, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
@@
Next, let $w_{n}=u_{s}e_{5}\in{\mathcal{A}}(e_{5})$ for some $u_{s}\in I(G)^{N-1}$. First observe that $(a,x)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:u_{s}e_{5})=(a)$ and $(u_{s}e_{3}:u_{s}e_{5})=(x)$. The case $w_{i}\in{\mathcal{A}}(e_{5})$ follows from the induction hypothesis as in the previous cases. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})$. Note that $w_{i}$ is divisible by $a$ or $x$ and $w_{n}$ is not divisible by $a$. Furthermore, $w_{n}$ is not divisible by $x$. Otherwise, if $x\mid u_{s}$, then $u_{s}=w^{\prime}e_{7}$ for a generator $w^{\prime}$ of $I(G)^{N-2}$ and $w_{n}=w^{\prime}e_{7}e_{5}=w^{\prime}e_{6}e_{3}$ which implies that $e_{3}\shortmid w_{n}$, a contradiction. So, $w_{i}:w_{n}\in(a,x)$. Hence, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
@@
Let $w_{n}=u_{s}e_{6}\in{\mathcal{A}}(e_{6})$ for some $u_{s}\in I(G)^{N-1}$. Observe that $(a)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{4}:u_{s}e_{6})=(a)$. The case $w_{i}\in{\mathcal{A}}(e_{6})$ follows from the induction hypothesis as in the previous cases. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})\cup{\mathcal{A}}(e_{5})$. Note that $w_{i}$ is divisible by $a$ or $b$ while $a\nshortmid w_{n}$ and $b\nshortmid w_{n}$. This implies that $w_{i}:w_{n}\in(a,b)$.
@@
Observe that $u_{s}=e_{6}^{t_{6}}e_{7}^{t_{7}}e_{8}^{t_{8}}$ since $u_{s}$ must be divisible by $z$. If $t_{8}\geq 1$, then $u_{s}=w^{\prime}e_{8}$ for a generator $w^{\prime}\in I(G)^{N-2}$ and $(w^{\prime}e_{6}e_{5}:w^{\prime}e_{6}e_{8})=(b)$. Similarly, if $t_{7}\geq 1$, then $u_{s}=w^{\prime}e_{7}$ for a generator $w^{\prime}\in I(G)^{q-2}$ and $(w^{\prime}e_{6}e_{3}:w^{\prime}e_{6}e_{7})=(b)$. So, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in these cases. For the remaining case, we have $w_{n}=e_{6}^{q}$. Notice that, for each $w_{i}$, we have $w_{i}:w_{n}\in(a,x,q)$ since $\{a,x,q\}$ is a vertex cover of $G\setminus\{e_{6}\}$. Furthermore, note that the generators $w_{j}=e_{6}^{N-1}e_{7}$ and $w_{k}=e_{6}^{N-1}e_{8}$ come before $w_{n}$ in ${\mathcal{A}}(e_{6})$ from the construction of ${\mathcal{N}}$. In addition, $(w_{j}:w_{n})=(x)$ and $(w_{j}:w_{n})=(q)$. Thus, the claim holds.
@@
For the final two cases, if $w_{n}=u_{s}e_{7}\in{\mathcal{A}}(e_{7})$ then none of $a,b,p$ divide $u_{s},$ but every element of ${\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})\cup{\mathcal{A}}(e_{5})\cup{\mathcal{A}}(e_{6})$ is divisible by one of these three variables. Since $(u_{s}e_{2}:u_{s}e_{7})=(a)$ and $(u_{s}e_{3}:u_{s}e_{7})=(b)$ and $(u_{s}e_{6}:u_{s}e_{7})=(p)$, we are done. For the last case, we have $w_{n}=e_{8}^{q}$. Note that $(e_{8}^{q-1}e_{5}:e_{8}^{q})=(b),~{}~{}(e_{8}^{N-1}e_{6}:e_{8}^{N})=(p)$ and $(e_{8}^{N-1}e_{7}:e_{8}^{N})=(x)$. Since $\{b,p,x\}$ is a vertex cover of $G\setminus\{e_{8}\}$, we have $w_{i}:w_{n}\in(b,p,x)$ for each $w_{i}\neq w_{n}$. Therefore, the claim holds.
We now consider a different class of gapfree graphs built from adding a vertex to a pentagon.
@@
Let $G$ be the finite gapfree graph given in Figure 3. Then $I(G)^{N}$ has linear quotients for all $N\geq 2$.
@@
Assume $N\geq 3$. Suppose that $I(G)^{N-1}$ has linear quotients with the ordering ${\mathcal{N}}$: $u_{1}>\cdots>u_{r}$ where $u_{1}=e_{1}^{N-1}$ and $u_{r}=e_{8}^{N-1}$. Let ${\mathcal{N}}^{(1)}:~{}~{}w_{1}>\cdots>w_{t}$ denote the efficient ordering of $I(G)^{N}$ coming from ${\mathcal{N}}$ as in Conjecture 3.1. As before, use the following notation to refer to elements in ${\mathcal{N}}^{(1)}$
${\mathcal{U}}e_{1}={\mathcal{A}}(e_{1}),{\mathcal{A}}(e_{2}),{\mathcal{A}}(e_{3}),{\mathcal{A}}(e_{4}),{\mathcal{A}}(e_{5}),{\mathcal{A}}(e_{6}),{\mathcal{A}}(e_{7}),{\mathcal{A}}(e_{8}),{\mathcal{A}}(e_{9}),$

### v2

@@ 5. Pentagon together with one vertex
The edge ideal $I(G)$ has $8$ generators and, since $e_{2}e_{6}=e_{4}e_{7}$ and $e_{3}e_{8}=e_{5}e_{7}$, it follows that $I(G)^{2}$ has $\binom{8+1}{2}-2=36-2=34$ generators. Let ${\mathcal{N}}^{(2)}$ denote the following order on these generators:
@@
It is now a direct computation to verify that with the order ${\mathcal{N}}^{(2)}$, $I(G)^{2}$ has linear quotients.
Using the ordering ${\mathcal{N}}^{(2)}$ of the generators of $I(G)^{2}$ as above, one can inductively show that all powers of $I(G)$ have linear quotients. The following notation will be useful in the proof of Proposition 5.2.
Notation 5.1.
Let $G$ be a finite graph and $w$ be one of the generators of $I(G)^{s}$ for $s\geq 1$. The notation $e_{j}\shortmid w$ means that there exists a factorization $w=e_{i_{1}}\cdots e_{i_{N}}$ with $e_{j}=e_{i_{k}}$ for some $1\leq k\leq s$.
@@
Let $G$ be the finite graph in Figure $2$. Then all powers $I(G)^{s}$ with $s\geq 2$ have linear quotients.
@@
We prove that Conjecture 4.1 is true for $G$ with $q_{0}=2$. Assume $s\geq 3$ and suppose that $I(G)^{s-1}$ has linear quotients with the ordering ${\mathcal{N}}^{(s-1)}$: $u_{1}>\cdots>u_{r}$ where $u_{1}=e_{1}^{s-1}$ and $u_{r}=e_{8}^{s-1}$. Let ${\mathcal{N}}^{(s)}:~{}~{}w_{1}>\cdots>w_{t}$ denote the efficient ordering of $I(G)^{s}$ constructed from ${\mathcal{N}}^{(s-1)}$ as in Conjecture 4.1. For ease of notation, use the following notation to refer to elements in ${\mathcal{N}}^{(s)}$
${\mathcal{A}}(e_{1}),{\mathcal{A}}(e_{4}),{\mathcal{A}}(e_{2}),{\mathcal{A}}(e_{3}),{\mathcal{A}}(e_{5}),{\mathcal{A}}(e_{6}),{\mathcal{A}}(e_{7}),{\mathcal{A}}(e_{8})$
where $w_{k}\in{\mathcal{A}}(e_{i})$ means $e_{i}\shortmid w_{k}$ and $e_{j}\nshortmid w_{k}$ for for $j$ such that $e_{j}^{s-1}>e_{i}^{s-1}$ in ${\mathcal{N}}$.
@@
CASE 1: Let $w_{n}=u_{\ell}e_{4}\in{\mathcal{A}}(e_{4})$ for $\ell\leq r$. First, observe that $(b)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{\ell}e_{1}:u_{\ell}e_{4})=(b)$ for $u_{\ell}e_{1}\in{\mathcal{A}}(e_{1})$. Let $w_{i}=u_{t}e_{4}\in{\mathcal{A}}(e_{4})$ for $t<\ell$. By the induction hypothesis, there exists $u_{l}$ with $l<\ell$ such that $u_{l}:u_{\ell}$ is a variable dividing $u_{t}:u_{\ell}$. Let $w_{j}=u_{l}e_{4}$ and notice that $j<n$. Then, $w_{j}:w_{n}$ is a variable that divides $w_{i}:w_{n}$, as desired. Note that $w_{j}\in{\mathcal{A}}(e_{1})$ is possible.
Next, let $w_{i}\in{\mathcal{A}}(e_{1})$. Recall that $b\nmid e_{4}$. If $b\nmid u_{\ell}$, then $w_{i}:w_{n}\in(b)$ since every element of ${\mathcal{A}}(e_{1})$ is divisible by $b$. Then, $(u_{1}e_{1},\ldots,u_{r}e_{1}):w_{n}=(b)$. Hence, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case. Suppose $b\mid u_{\ell}$. Then either $e_{3}\shortmid u_{\ell}$ or $e_{5}\shortmid u_{\ell}$. If $e_{3}\shortmid u_{\ell}$, then $u_{\ell}=w^{\prime}e_{3}$ for a generator $w^{\prime}\in I(G)^{s-2}$. Consider the elements $w_{j}=w^{\prime}e_{4}e_{1}$ and $w_{k}=w^{\prime}e_{7}e_{1}$ in ${\mathcal{A}}(e_{1})$. Note that $(w_{j}:w_{n})=(a)$ and $(w_{k}:w_{n})=(z)$. If $e_{5}\shortmid u_{\ell}$, then $u_{\ell}=w^{\prime\prime}e_{5}$ for a generators $w^{\prime\prime}\in I(G)^{s-2}$. Similarly, consider $w_{j}=w^{\prime\prime}e_{4}e_{1}$ and $w_{k}=w^{\prime\prime}e_{8}e_{1}$ in ${\mathcal{A}}(e_{1})$ where $(w_{j}:w_{n})=(a)$ and $(w_{k}:w_{n})=(z)$. So, in either of these scenarios, we have $(a,b,z)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}.$ Let $w_{i}=e_{1}^{t_{1}}\cdots e_{8}^{t_{8}}$ where $t_{1}>0$ and $w_{n}=e_{2}^{t_{2}^{\prime}}\cdots e_{8}^{t_{8}^{\prime}}$ since $e_{1}\nshortmid w_{n}$. If $w_{i}:w_{n}\notin(a,b,z)$, then examining the degrees of $a,b,z$ in $w_{i}$ and $w_{n}$, respectively, results in
@@
CASE 3: Let $w_{n}=u_{\ell}e_{2}\in{\mathcal{A}}(e_{2})$ for some $\ell\leq r$. Note that $e_{6}\nshortmid w_{n}$ because $e_{6}e_{2}=e_{7}e_{4}$ and $e_{4}\nshortmid e_{n}$. Next, observe that $(b,p)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{\ell}e_{1}:u_{\ell}e_{2})=(b)$ and $(u_{\ell}e_{4}:u_{\ell}e_{2})=(p)$. Let $w_{i}\in{\mathcal{A}}(e_{2})$. As in the previous case, by the induction hypothesis, there exists $w_{j}$ for $j<n$ such that $w_{j}:w_{n}$ is a variable that divides $w_{i}:w_{n}$, as desired. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})$. Observe that $p\nmid u_{\ell}$; otherwise, $e_{4}\shortmid u_{\ell}$ or $e_{6}\shortmid u_{\ell}$ but neither of these cases are possible since $w_{n}\in{\mathcal{A}}(e_{2})$. If $b\nmid u_{\ell}$, then $(w_{i}:w_{n})\subseteq(b,p)$ which implies that $(w_{1},\ldots,w_{n-1}):w_{n}$ is generated by variables. So, claim holds for this case. Suppose $b\mid u_{\ell}.$ Then either $e_{3}\shortmid u_{\ell}$ or $e_{5}\shortmid u_{\ell}$ since $e_{1}\nshortmid u_{\ell}.$ If $e_{3}\shortmid u_{\ell}$, then $u_{\ell}=w^{\prime}e_{3}=w^{\prime}bx$ for some $w^{\prime}\in I(G)^{s-2}$. Consider the elements $w_{j}=w^{\prime}e_{2}e_{1}$ and $w_{k}=w^{\prime}e_{7}e_{1}$ in ${\mathcal{A}}(e_{1})$ where $(w_{j}:w_{n})=(a)$ and $(w_{k}:w_{n})=(z)$. So, $(a,b,p,z)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}.$
@@
Assuming $e_{5}\shortmid u_{\ell}$ also yields to the above containment. Let $w_{i}=e_{1}^{t_{1}}\cdots e_{8}^{t_{8}}$ where $t_{1}>0$ or $t_{4}>0$ and $w_{n}=e_{2}^{t_{2}^{\prime}}e_{3}^{t_{3}^{\prime}}e_{5}^{t_{5}^{\prime}}e_{t}^{t_{7}^{\prime}}e_{8}^{t_{8}^{\prime}}$ since $e_{1},e_{4},e_{6}$ do not divide $w_{n}=u_{\ell}e_{2}.$ If $w_{i}:w_{n}\notin(a,b,p,z)$, then examining degrees of $a,b,p,z$ in $w_{i}$ and $w_{n}$ results with $t_{1}+t_{4}+t_{6}\leq 0$, a contradiction. So, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
@@
CASE 4: Now, let $w_{n}=u_{\ell}e_{3}\in{\mathcal{A}}(e_{3})$ for some $u_{\ell}\in I(G)^{s-1}$. First observe that $(a)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{\ell}e_{1}:u_{\ell}e_{3})=(a)$. The case $w_{i}\in{\mathcal{A}}(e_{3})$ follows from the induction hypothesis as in the previous cases. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})\cup{\mathcal{A}}(e_{2})$. Note that this collection contains all generators of $I(G)^{s}$ that are divisible by $a$ since $e_{1},e_{2}$ and $e_{4}$ are the only edges incident to $a$ and $a\nmid w_{n}$. This means $w_{i}:w_{n}\in(a)$. Hence, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
@@
Next, let $w_{n}=u_{\ell}e_{5}\in{\mathcal{A}}(e_{5})$ for some $u_{\ell}\in I(G)^{s-1}$. First observe that $(a,x)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{\ell}e_{1}:u_{\ell}e_{5})=(a)$ and $(u_{\ell}e_{3}:u_{\ell}e_{5})=(x)$. The case $w_{i}\in{\mathcal{A}}(e_{5})$ follows from the induction hypothesis as in the previous cases. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})$. Note that $w_{i}$ is divisible by $a$ or $x$ and $w_{n}$ is not divisible by $a$. Furthermore, $w_{n}$ is not divisible by $x$. Otherwise, if $x\mid u_{\ell}$, then $u_{\ell}=w^{\prime}e_{7}$ for a generator $w^{\prime}$ of $I(G)^{s-2}$ and $w_{n}=w^{\prime}e_{7}e_{5}=w^{\prime}e_{6}e_{3}$ which implies that $e_{3}\shortmid w_{n}$, a contradiction. So, $w_{i}:w_{n}\in(a,x)$. Hence, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
@@
CASE 5: Let $w_{n}=u_{\ell}e_{6}\in{\mathcal{A}}(e_{6})$ for some $u_{\ell}\in I(G)^{s-1}$. Observe that $(a)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{\ell}e_{4}:u_{\ell}e_{6})=(a)$. The case $w_{i}\in{\mathcal{A}}(e_{6})$ follows from the induction hypothesis as in the previous cases. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})\cup{\mathcal{A}}(e_{5})$. Note that $w_{i}$ is divisible by $a$ or $b$ while $a\nshortmid w_{n}$ and $b\nshortmid w_{n}$. This implies that $w_{i}:w_{n}\in(a,b)$.
@@
Observe that $u_{\ell}=e_{6}^{t_{6}}e_{7}^{t_{7}}e_{8}^{t_{8}}$ since $u_{\ell}$ must be divisible by $z$. If $t_{8}\geq 1$, then $u_{\ell}=w^{\prime}e_{8}$ for a generator $w^{\prime}\in I(G)^{s-2}$ and $(w^{\prime}e_{6}e_{5}:w^{\prime}e_{6}e_{8})=(b)$. Similarly, if $t_{7}\geq 1$, then $u_{\ell}=w^{\prime}e_{7}$ for a generator $w^{\prime}\in I(G)^{s-2}$ and $(w^{\prime}e_{6}e_{3}:w^{\prime}e_{6}e_{7})=(b)$. So, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in these cases. For the remaining case, we have $w_{n}=e_{6}^{s}$. Notice that, for each $w_{i}$, we have $w_{i}:w_{n}\in(a,x,q)$ since $\{a,x,q\}$ is a vertex cover of $G\setminus\{e_{6}\}$. Furthermore, note that the generators $w_{j}=e_{6}^{s-1}e_{7}$ and $w_{k}=e_{6}^{s-1}e_{8}$ come before $w_{n}$ in ${\mathcal{A}}(e_{6})$ from the construction of ${\mathcal{N}}^{(s)}$. In addition, $(w_{j}:w_{n})=(x)$ and $(w_{k}:w_{n})=(q)$. Thus, the claim holds.
@@
CASE 6: For the final two cases, if $w_{n}=u_{\ell}e_{7}\in{\mathcal{A}}(e_{7})$ then none of $a,b,p$ divide $u_{\ell},$ but every element of ${\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{4})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})\cup{\mathcal{A}}(e_{5})\cup{\mathcal{A}}(e_{6})$ is divisible by one of these three variables. Since $(u_{\ell}e_{2}:u_{\ell}e_{7})=(a)$ and $(u_{\ell}e_{3}:u_{\ell}e_{7})=(b)$ and $(u_{\ell}e_{6}:u_{\ell}e_{7})=(p)$, we are done. For the last case, we have $w_{n}=e_{8}^{s}$. Note that $(e_{8}^{s-1}e_{5}:e_{8}^{s})=(b),~{}~{}(e_{8}^{s-1}e_{6}:e_{8}^{s})=(p)$ and $(e_{8}^{s-1}e_{7}:e_{8}^{s})=(x)$. Since $\{b,p,x\}$ is a vertex cover of $G\setminus\{e_{8}\}$, we have $w_{i}:w_{n}\in(b,p,x)$ for each $w_{i}\neq w_{n}$. Therefore, the claim holds. ∎
As an example of how 3.4 can be applied, consider the class of graphs formed by replacing the central vertex $x$ of Figure 3 by a complete graph $K_{n}$, with each vertex of $K_{n}$ connected to $a,b,z$. We denote such a graph by $(C_{5},K_{n})$ and label the vertices of $K_{n}$ by $x_{1},\ldots,x_{n}$. The case $n=4$ is shown below.
$z$$p$$q$$x_{1}$$x_{2}$$x_{3}$$x_{4}$$a$$b$
Figure 4. $(C_{5},K_{4})$
Example 5.3.
Let $I$ be the edge ideal of the graph $(C_{5},K_{n})$ for some $n\geq 1$. Then $I^{s}$ has linear quotients for all $s\geq 2$. The case $n=1$ holds by 5.2. For $n>1$ the result holds via induction using Proposition 2.3.
Notice that the above argument cannot extend to replace any vertex of the exterior pentagon with an edge. In particular, this would result in a finite graph that is not gapfree. However, the technique can be applied to the central vertex in Figure 5, for instance, which is the base case for a second class of gapfree graphs built from adding a vertex to a pentagon.
@@
Let $G$ be the finite gapfree graph given in Figure 5. Then $I(G)^{s}$ has linear quotients for all $s\geq 2$.
@@
Assume $s\geq 3$. Suppose that $I(G)^{s-1}$ has linear quotients with the ordering ${\mathcal{N}}^{(s-1)}$: $u_{1}>\cdots>u_{r}$ where $u_{1}=e_{1}^{s-1}$ and $u_{r}=e_{9}^{s-1}$. Let ${\mathcal{N}}^{(s)}:~{}~{}w_{1}>\cdots>w_{t}$ denote the efficient ordering of $I(G)^{s}$ constructed from ${\mathcal{N}}^{(s-1)}$ as in Conjecture 4.1. As before, use the following notation to refer to elements in ${\mathcal{N}}^{(s)}$
${\mathcal{A}}(e_{1}),{\mathcal{A}}(e_{2}),{\mathcal{A}}(e_{3}),{\mathcal{A}}(e_{4}),{\mathcal{A}}(e_{5}),{\mathcal{A}}(e_{6}),{\mathcal{A}}(e_{7}),{\mathcal{A}}(e_{8}),{\mathcal{A}}(e_{9}),$